solve using the elimination method.
8x-5y=-9
3x+5y=-2

Answer by mananth(16946)

8x-5y=-9
3x+5y=-2
Add the two equations
8x-5y+3x+5y=-9-2
11x=-11
x= -11/11
x=-1
8x-5y=-9
-8-5y=-9
-5y=-9+8
y=-1/5
Ananth